Gd3+ has seven unpaired electrons, find its spin only magnetic moment.


1
Expert's answer
2021-04-19T04:34:41-0400

Magnetic Moment ="\sqrt(n(n+2)"


Magnetic Moment ="\sqrt (7(7+2)=\sqrt 63"


Magnetic Moment ="7.9BM"